OpenStudy (thickgirl_angel):

LAW ENFORCEMENT: A police accident investigator can use the formula S=2√ 5L (5L is under the radical) to estimate the speed s of a car in miles per hour based on the length l in feet of the skid marks it left. How fast was a car traveling that left skid marks 105 feet long?

OpenStudy (phi):

\[ S = 2 \sqrt{5L} \] when L is 105 feet we do \[ S = 2 \sqrt{5\cdot 105}\]

OpenStudy (thickgirl_angel):

45.825